I told Koji that I wanted to play with the common mode servo this evening, and he pointed out that we only get the signals after the digital demod phase angle in the digital system (obviously). So, if I want to use either REFL11 or REFL55 for my CARM signal, I want to do something in analog-land so that my digital demod phase is close to 0 or 90.
While we had the PRFPMI locked (with CARM offset of 2 or 3 nm), we set the demod phases of REFL11 and REFL55 to minimize a CARM line in the Q-phase. This gave us -34 degrees for REFL11, and -75 degrees for REFL55.
We calculated that about 1 degree of phase shift is about 1/(2 * pi * freq), or about 1.4e-8 seconds of delay for 11MHz. We took the speed of light in the cables to be about 2/3*c, so 1.4e-8 * 2e8 = 2.9 meters per degree for 11MHz. Since REFL11 was 34 degrees from 0, we estimate that we need to add about 98 meters of cable to the REFL11 signal path. The same calculation for 55 MHz, but with a 15 degree shift required, gives 8.8 meters of cable to be added to the REFL55 signal path.
I connected up some long BNC cables, and inserted them between the heliax breakout board on the LSC rack, and the respective PD inputs of the REFL11 and REFL55 demod boards. I used (45 meters + 45 meters + a little bit) for REFL11, and used about 9 meters for REFL55.
When we relocked the PRFPMI, and redid the phasing, we were very close to zero for both REFL11 and REFL55! REFL11's digital demod phase is now +1 degree, and REFL55's digital demod phase is -5 degrees.
We changed the input of the CM servo board from POY (which Den and Koji had been using in December - see elog 9500) to REFL11 I MON.
Q locked the FPMI (separate reply elog), and then we tried engaging the CM analog servo. We were not successful.
These settings were mostly copied from elog 9500, so they are almost surely not correct.
CM servo screen: In1 gain = 31dB, switch on, offset = -2.7V, boost off, super boosts off, option=disable, 79:1.6k switch disabled, polarity minus, option disable, AO gain=8dB, limiter enable.
For the slow path, CM_SLOW -> MC LSC servo had a +1 in the input matrix.
CM filters in the AUX_ERR screen: FM1 (unwhite) on, all others off, gain = 2.6.
MC servo filters: FM7, FM10 on, all others off (no triggered filter modules). Gain = 0 initially.
MC servo board AO path disabled initially, G=-32dB initially.
Once Q had the FPMI locked, I tried increasing just the CM analog gain (by enabling the AO path on the MC board, and increasing the gain). Doing this, I lost lock at -3 dB.
I then tried again, this time alternating increasing the analog gain, and increasing the MC LSC servo gain. I got up to 3e-3 for the MC digital gain, and -7 dB for the analog gain before we lost lock again.
We have determined that we should probably try just locking one of the arms with POX or POY, as Den and Koji did, to get a feel for how the system works.